OpenStudy (anonymous):

What is: 8/5 divided by 40/12

OpenStudy (anonymous):

this expression can be written as 8*12/5*40=96/200=0.48

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

8/5 / 40/12 = 1/3

OpenStudy (texaschic101):

when dividing fractions, " flip " the number that you are dividing by, then multiply 8/5 divided by 40/12 = 8/5 * 12/40 = 96/200 which reduces to 12/25

OpenStudy (anonymous):

Oh...

OpenStudy (texaschic101):

make sure you flip the number you are dividing by, and not the other number.